# StreamLTS: Query-based Temporal-Spatial LiDAR Fusion for Cooperative Object Detection

### v1

@@ Abstract
Cooperative perception via communication among intelligent traffic agents has great potential to improve the safety of autonomous driving. However, limited communication bandwidth, localization errors and asynchronized capturing time of sensor data, all introduce difficulties to the data fusion of different agents. To some extend, previous works have attempted to reduce the shared data size, mitigate the spatial feature misalignment caused by localization errors and communication delay. However, none of them have considered the asynchronized sensor ticking times, which can lead to dynamic object misplacement of more than one meter during data fusion. In this work, we propose Time-Aligned COoperative Object Detection (TA-COOD), for which we adapt widely used dataset OPV2V and DairV2X with considering asynchronous LiDAR sensor ticking times and build an efficient fully sparse framework with modeling the temporal information of individual objects with query-based techniques. The experiment results confirmed the superior efficiency of our fully sparse framework compared to the state-of-the-art dense models. More importantly, they show that the point-wise observation timestamps of the dynamic objects are crucial for accurate modeling the object temporal context and the predictability of their time-related locations.
@@ 1 Introduction
In recent years, cooperative perception with sharing information among Intelligent Agents (IAs), such as Connected Autonomous Vehicles (CAVs) and Connected Infrastructures (CIs), has proven to be beneficial for improving the perception accuracy and road safety[4, 26, 34, 22, 30]. Different from the ego-perception system that only observes the driving environment from the ego vehicle’s view, cooperative perception collect information from the sensors mounted on different IAs. These sensors typically have better spatial arrangement, can efficiently mitigate the occlusions and increase the perception distance of the ego vehicle. This ensures safer driving and more optimized route planning of AVs. However, sharing and fusing information from different IAs introduces several challenges, including the limited communication bandwidth and information misalignment introduced by localization errors and asynchronized capturing time of sensor data.
@@
Previous works have sought to minimize bandwidth usage by compressing the learned deep Bird’s Eye View (BEV) features [22, 4, 26] or selecting deep keypoint features [34], rectify localization errors through learning-based techniques [22, 25] or analytic algorithms [34, 33], and synchronize communication delays with the attention mechanism [24] or explicitly predicting the future features before data fusion [31]. Nonetheless, none have considered the asynchronized sensor ticking time.
@@
More specifically, previous benchmarks [26, 30, 24, 31] assume that the sensors have synchronized global ticking time. In reality, these sensors might have asynchronous ticking time in each aligned frame, leading to inhomogeneous observation time of each object in the scenario. This effect can be amplified when the rolling shutter effect (e.g. mechanic LiDARs) exists. At the sensor capturing frequency of 10Hz, these time differences ($0~{}\text{to}~{}100ms$), that are in the same magnitude as the communication delay, can lead to large spatial displacement at data fusion, especially when the CAVs have high speeds. Therefore, they should not be ignored. As the two example scenes shown in Fig. 1, we leverage the timestamps of each point (illustrated with differently colored points) in the point cloud to achieve globally Time-Aligned COoperative Object Detection (TA-COOD) that predicts object bounding boxes (BBoxes) for the unified global timestamp (green boxes) in stead of detecting the timely asynchronous local bounding boxes (yellow boxes).
@@
To model more robust temporal features, we utilize the information from the previous frames. This is very computational-demanding as the data from several IAs should be processed. Most of the previous works [26, 22, 28, 25, 8] use dense convolutions or transformers to process BEV feature maps, that consume a lot of GPU memory and are hard to extend the model to sequential data on limited computational resources. In contrast, the potential of fully sparse network structures has rarely been researched. In this work, we built a fully sparse framework that can be trained on a single Nvidia GTX 3090 with sufficient batch size (e.g. 4) and processing sequential LiDAR streaming data of more than 4 frames and up to 7 CAVs.
@@ 2.1 Cooperative Object Detection
However, semi-processed data can be in any format. For instance, Fcooper [4] experimented with both voxel feature and spatial deep feature map fusion, found that voxel feature fusion has better performance as it keeps more details. Instead, [26] encodes the point clouds into point-pillars [11] with fully-connected layers and uses dense convolutions to generate feature maps. These feature maps from different IAs are then fused with an attention module instead of max-out used in Fcooper. These learnable encoding and fusion modules ensure that the most important information is selected and shared to achieve better performance. To further increase the detection accuracy while spatial misalignment between the feature maps of IAs exists, SCOPE [27] introduced the pyramid LSTM [12] to reason about temporal and spatial information from sequential point clouds.
While these works all take feature maps as their fusion strategy, [10] and [35] utilize the learned criterion, such as confidence and uncertainty, to select the most relevant information and reduce the data size for sharing. Similarly, [34] adopt the sparsity of point clouds and learn the most important object keypoints for sharing to further reduce the bandwidth requirement. They all achieve good performances on the OPV2V [26] dataset with significantly less data sharing.
@@ 2.2 Query-based Temporal Data Fusion
Based on transformer, DETR [2] first introduced the query-based object detection for 2D images. Since it is anchor-free and needs no post-processing, it has been widely used and was extended to 3D object detection [23, 15, 16]. More importantly, the object queries provide great convenience to efficiently interact with features of any modality. For instance, by integrating the 3D geometry information into the image memory features, PETR [15] built a simple yet effective framework to let 3D queries directly interact with the 2D features in image space and obtained superior 3D object detection performance. By propagating the learned object queries to the next frames and model the temporal context for the current frame, QueryProp [9] improved the efficiency and accuracy for the video object detection. With only images as input, StreamPETR [21] achieved on-par 3D object detection performance to the LiDAR-track benchmark on the NuScenes [1] dataset. Further more, the concept of object query propagation has been applied to object tracking tasks. MOTR [36] and TrackFormer [18] use object queries to achieve temporal modeling and associate the objects between frames in an end-to-end manner, which is simple yet effective on tracking. Inspired by these works, we model the temporal information of observations from different IAs with the selected potential object queries, and share the learned queries for TA-COOD to save communication bandwidth.
@@ 2.3 Datasets for Collective Perception
Recently, the real datasets DAIR-V2X [30] and V2V4Real [24] were made accessible. Both are only configured with two agents, DAIR-V2X with one CAV and one connected infrastructure, V2V4Real with two CAVs. Because of the dynamics and asynchrony of sensors, there might be spatial misalignment between dynamic objects observed by the two agents. To generate unified ground-truth BBoxes for cooperative object perception, both DAIR-V2X and V2V4Real take ego-vehicle’s annotation as the ground-truth if the annotations from two agents spatially overlap. However, at the blind spot of the ego-vehicle, the annotations of the cooperative agent are taken. In this way, the final generated bounding box might contain errors and does not reflect the true location of the encapsulated object at the given timestamp. This might also negatively influence the temporal predictability of sequential models. To better preserve the detailed temporal information and its relation to the accurate location of the objects in the scenario, we proposed to generate global time-aligned ground-truth BBoxes for the TA-COOD task.
@@ 3.2 Overall Framework
Figure 3: Left: Overall framework for TA-COOD. Local Data Process encodes local point cloud features and fuse the temporal context into the object queries. Data Fusion Process fuses the spatial object query features from the $N$ IAs. Right: MTA alignment. Motion-Time-Aware embeds the query positions ($Q_{p}$) and Motion-Aware embeds the query context $Q_{c}$ and initial queries ($tgt$).
@@ 3.3 Module Configuration
[…]
@@ 4.1 Datasets
OPV2Vt (Fig. 1(a)): The orignal OPV2V dataset is generated by the simulator CARLA [6]. It contains about 12K frames of data in 73 scenes. We interpolate the meta data of OPV2V to generate ten intermediate locations for each object and sensor inside one frame, which leads to ten sub-frames (0.01s) in each scan frame (0.1s) of LiDARs. We then replay all sub-frames to generate data with more fine-grained dynamics in each scan frame. Since these sub-frames are synchronized, we randomly drop the data of several beginning sub-frames (1 to 5) for each scan sequence and joint every consecutive 10 sub-frames as a full LiDAR scan to simulate the ticking time asynchrony. For each jointed frame data, we use the BBoxes of the last sub-frame as ground-truth. One frame of data is shown in Fig. 1(a). The evaluation detection range is set to $[-140.8,140.8]m$, $[-38.4,38.4]m$, and $[-3.0,1.0]m$ for x-, y- and z-axis, respectively.
DairV2Xt (Fig. 1(b)): The real dataset DairV2X is captured with one CAV and one CI at several intersections. To generate the globally time-aligned BBoxes, we first refine the point cloud alignment between the CAV and the CI to decouple the bounding box misalignment caused by localization errors and asynchronized capturing times. In each sequence, we select the frame where CAV is closest to CI as the global reference coordinate $O_{ref}$, and then register sequentially the point clouds of CAV of other frames to $O_{ref}$. During registration, dynamic points are moved via the BBox annotations. Based on the registration, we obtain a global merged and down-sampled point cloud, to which the CI point clouds are then registered to obtain the relative location to $O_{ref}$. With the corrected sensor poses in each frame, the annotated objects are transformed from the local sensor coordinate to the globally aligned coordinate $O_{ref}$ and frame-wisely aligned to the object trajectories. By interpolation, we generate the ground-truth BBoxes that are aligned to the scan end $t_{aligned}$ in each frame. The detection range at y- and z-axis is the same as OPV2Vt. The range for x-axis is set to $[-102.4,102.4]m$.

### v2

@@ Abstract
Cooperative perception via communication among intelligent traffic agents has great potential to improve the safety of autonomous driving. However, limited communication bandwidth, localization errors and asynchronized capturing time of sensor data, all introduce difficulties to the data fusion of different agents. To some extend, previous works have attempted to reduce the shared data size, mitigate the spatial feature misalignment caused by localization errors and communication delay. However, none of them have considered the asynchronized sensor ticking times, which can lead to dynamic object misplacement of more than one meter during data fusion. In this work, we propose Time-Aligned COoperative Object Detection (TA-COOD), for which we adapt widely used dataset OPV2V and DairV2X with considering asynchronous LiDAR sensor ticking times and build an efficient fully sparse framework with modeling the temporal information of individual objects with query-based techniques. The experiment results confirmed the superior efficiency of our fully sparse framework compared to the state-of-the-art dense models. More importantly, they show that the point-wise observation timestamps of the dynamic objects are crucial for accurate modeling the object temporal context and the predictability of their time-related locations. The official code is available at https://github.com/YuanYunshuang/CoSense3D.
@@ 1 Introduction
In recent years, cooperative perception with sharing information among Intelligent Agents (IAs), such as Connected Autonomous Vehicles (CAVs) and Connected Infrastructures (CIs), has proven to be beneficial for improving the perception accuracy[4, 28, 36, 24, 32]. Different from the ego-perception system that only observes the driving environment from the ego-view, cooperative perception collect information from the sensors mounted on different IAs. These sensors typically have better spatial arrangement, can efficiently mitigate the occlusions and increase the perception distance of the ego vehicle. This ensures safer driving and more optimized route planning of AVs. However, sharing and fusing information from different IAs introduces several challenges, including the limited communication bandwidth and information misalignment introduced by localization errors and asynchronized capturing time of sensor data.
@@
Previous works have sought to minimize bandwidth usage by compressing the learned deep Bird’s Eye View (BEV) features [24, 4, 28, 5, 8] or selecting deep keypoint features [36], rectify localization errors through learning-based techniques [24, 27] or analytic algorithms [36, 35], and synchronize communication delays with the attention mechanism [26] or explicitly predicting the future features before data fusion [33]. Nonetheless, none considered the asynchronized sensor ticking time.
@@
More specifically, previous benchmarks [28, 32, 26, 33] assume that the sensors have synchronized ticking time in each aligned frame. In reality, sensors tick asynchronously, leading to inhomogeneous observation time of each object in the scenario. This effect can be amplified when the rolling shutter effect (e.g. mechanic LiDARs) exists. At the sensor capturing frequency of 10Hz, these time differences ($0~{}\text{to}~{}100ms$), that are in the same magnitude as the communication delay, can lead to large spatial displacement at data fusion, especially when the CAVs have high speeds. Therefore, they should not be ignored. As the two example scenes shown in Fig. 1, we leverage the timestamps of each point (illustrated with differently colored points) in the point cloud to achieve globally Time-Aligned COoperative Object Detection (TA-COOD) that predicts object bounding boxes (BBoxes) for the unified global timestamp (green boxes) in stead of detecting the timely asynchronous local bounding boxes (yellow boxes).
@@
To model more robust temporal features, we utilize the information of previous frames. This is very computational-demanding as the data from several IAs should be processed. Most previous works [28, 24, 30, 27, 10] use dense convolutions or transformers to process BEV feature maps, that consume a lot of GPU memory and are hard to extend the model to sequential data on limited computational resources. In contrast, the potential of fully sparse network structures has rarely been researched. In this work, we built a fully sparse framework that can be trained on a single Nvidia GTX 3090 with sufficient batch size (e.g. 4) and processing LiDAR streaming data of more than 4 frames and up to 7 CAVs.
@@ 2.1 Cooperative Object Detection
However, semi-processed data can be in any format. For instance, Fcooper [4] experimented with both voxel feature and feature map fusion, found that voxel feature fusion has better performance as it keeps more details. Instead, [28] encodes the point clouds into point-pillars [13] with fully-connected layers and uses dense convolutions to generate feature maps of different IAs, and then fuse them with an attention module. These learnable encoding and fusion modules ensure that the most important information is selected and shared to achieve better performance. To further increase the detection accuracy despite spatial misalignment of feature maps, SCOPE [29] introduced the pyramid LSTM [14] to reason about temporal and spatial information from sequential point clouds.
While these works all take feature maps as their fusion strategy, [12] and [37] utilize the learned criterion, such as confidence and uncertainty, to select the most relevant information and reduce the data size for sharing. Similarly, [36] adopt the sparsity of point clouds and learn the most important object keypoints for sharing. To further reduce the bandwidth requirement, [5] selects the top ranking point features from the BEV feature map as object queries for further processing, sharing and fusing. They all achieve good performances with significantly less data sharing.
@@ 2.2 Query-based Temporal Data Fusion
Based on transformer, DETR [2] first introduced the query-based object detection for 2D images. Since it is anchor-free and needs no post-processing, it has been widely used and was extended to 3D object detection [25, 17, 18]. More importantly, the object queries provide great convenience to efficiently interact with features of any modality. For instance, by integrating the 3D geometry information into the image memory features, PETR [17] built a simple yet effective framework to let 3D queries directly interact with the 2D features in image space and obtained superior 3D object detection performance. By propagating the learned object queries to the next frames and model the temporal context for the current frame, QueryProp [11] improved the efficiency and accuracy for the video object detection. With only images as input, StreamPETR [23] achieved on-par 3D object detection performance to the LiDAR-track benchmark on the NuScenes [1] dataset. Further more, the concept of object query propagation has been applied to object tracking tasks. MOTR [38] and TrackFormer [20] use object queries to achieve temporal modeling and associate the objects between frames in an end-to-end manner, which is simple yet effective on tracking. QUEST [8] extended PETR to cooperative object detection without considering the temporal information. In this work, we model the temporal context of observations from different IAs with the selected potential object queries, and share the learned queries for TA-COOD to save communication bandwidth.
@@ 2.3 Datasets for Collective Perception
Recently, the real datasets DairV2X [32] and V2V4Real [26] were made accessible. Both are only configured with two agents, DairV2X with one CAV and one connected infrastructure, V2V4Real with two CAVs. Because of the dynamics and asynchrony of sensors, there might be spatial misalignment between dynamic objects observed by the two agents. To generate unified ground-truth BBoxes for cooperative object perception, both DairV2X and V2V4Real take ego-vehicle’s annotation as the ground-truth if the annotations from two agents spatially overlap. However, at the blind spot of the ego-vehicle, the annotations of the cooperative agent are taken. In this way, the final generated bounding box might contain errors and does not reflect the true location of the encapsulated object at the given timestamp. This might also negatively influence the temporal predictability of sequential models. To better preserve the detailed temporal information and its relation to the accurate location of the objects in the scenario, we proposed to generate global time-aligned ground-truth BBoxes for the TA-COOD task.
@@ 3.2 Overall Framework
Figure 3: Left: Overall framework for TA-COOD. Local Data Process encodes local point cloud features and fuse the temporal context into the object queries; Weights are shared across agents. Data Fusion Process fuses the spatial object query features from the $N$ IAs. Right: MTA alignment. Motion-Time-Aware embeds the query positions ($Q_{p}$) and Motion-Aware embeds the query context $Q_{c}$ and initial queries ($tgt$).
@@ 3.3 Module Configuration
After each iteration of TempFusion, the SpatialFusion module fuses the local object queries $Q_{i}^{t}$ from different IAs . As the example of three IAs ($A_{0},A_{1},A_{2}$) shown in the purple block of Fig. 4, the input query points $Q_{i}^{t}$ are firstly spatially aligned by taking the union of all query locations $P_{i}^{t}$ to obtain the same number of queries $\breve{P}_{i}^{t}$ at each IA. The query features at each padded location (gray cubes in Fig. 4) are initialized with zeros (Eq. 2). These queries are then used for the global TA-COOD.
@@ 4.1 Datasets
OPV2Vt (Fig. 1(a)): The orignal OPV2V dataset is simulated by CARLA [7]. It contains about 12K frames of data in 73 scenes. We interpolate the meta data of OPV2V to generate ten intermediate locations for each object and sensor inside one frame, which leads to ten sub-frames (0.01s) in each scan frame (0.1s). We then replay all sub-frames to generate data with more fine-grained dynamics. Since these sub-frames are synchronized, we randomly drop the data of several beginning sub-frames for each scan sequence and joint every consecutive 10 sub-frames as a full LiDAR scan to simulate the ticking time asynchrony. For each jointed frame data, we use the BBoxes of the last sub-frame as ground-truth. More details can be found in Appendix A. One frame of data is shown in Fig. 1(a). The evaluation detection range is set to $[-140.8,140.8]m$, $[-38.4,38.4]m$, and $[-3.0,1.0]m$ for x-, y- and z-axis, respectively.
DairV2Xt (Fig. 1(b)): The original DairV2X is captured with one CAV and one CI at intersections. To generate the globally time-aligned BBoxes, we first refine the point cloud alignment between the CAV and the CI to decouple the BBox misalignment caused by localization errors and asynchronized capturing times. In each sequence, we select the point cloud of the CAV in the sequence where the location of the CAV is closest to CI as the global reference $O_{ref}$, and then register sequentially the point clouds of both CAV and CI in other frames to $O_{ref}$. With the corrected sensor poses, the annotated objects are transformed from the local sensor coordinate to the globally aligned coordinate $O_{ref}$ and frame-wisely aligned to the object trajectories. By interpolation, we generate the ground-truth BBoxes that are aligned to the scan end $t_{aligned}$ in each frame. More details can be found in Appendix B. The detection range at y- and z-axis is the same as OPV2Vt. The range for x-axis is set to $[-102.4,102.4]m$.
@@ 6 Conclusion
Acknowledgements
This work is supported by the projects DFG RTC1931 SocialCars.
Appendix A: OPV2Vt Generation
The overall process of generating OPV2Vt dataset is described in Algorithm 1. This algorithm is applied to each scene of the OPV2V dataset. Note that the coordinate-related parameters $x,y,z,\theta$ are represented in the global coordinate system, where $\theta$ represents the yaw angle. The roll and pitch angles are set to 00. Additionally, note that $N_{s}<N_{v}$, and each sensor is installed on the roof of a vehicle in the set $\mathbf{v}$. The functions used in Algorithm 1 are defined as follows:
Interpolation$(v1,v2,n)$: Given the vehicle parameters $v1$ and $v2$ from two adjacent frames and the subframe index $1\leq n\leq 10$, we interpolate the coordinates $x,y,z$ and the orientation $\theta\in[0,2\pi)$ using Eq. 5. For the orientation, we additionally use Eq. 6 to resolve the direction flipping issue after interpolation.
$u$ $=u1+(u2-u1)\cdot\frac{n}{10}$ (5)
$\theta$ =$=\begin{cases}\theta-2\pi\quad\text{ if }\theta>\pi\\ \theta+2\pi\quad\text{ if }\theta<-\pi\end{cases}$ (6)
CarlaPlay$(\mathbf{v},\mathbf{s},t)$: Place all vehicles $\mathbf{v}$ and LiDAR sensors $\mathbf{s}$ in the CARLA simulator according to their parameters and generate the corresponding point cloud for the time interval $\delta t$. Note that we do not change the sensor recording frequency; each full scan requires $0.1$ seconds. Therefore, each subframe only generates data for $\delta t=0.01$ seconds, which corresponds to the scan points in a $36^{\circ}$ sector as shown in Fig. 1(a) (each sector is illustrated with a different color).
RandomInteger$(a,b)$: Select one integer in the range $[a,b]$.
Algorithm 1 OPV2Vt Generation
1:Input: OPV2V meta info $\mathcal{F}$
2:$\Delta t$triangleright Time interval between two adjacent frames
3:=$\mathcal{F}=\{(\mathbf{v}_{i},\mathbf{s}_{i},t_{i})|i=\{1,2,\dots,N_{f}\}\}$, $\triangleright$ Frame data $f_{i}$ at all timestamps $t_{i}$
4:$\mathbf{v}=\{v_{i}|i=\{1,2,\dots,N_{v}\}\}$ , $\triangleright$ Vehicle info in the i-th frame
5:$\mathbf{s}=\{s_{i}|i=\{1,2,\dots,N_{s}\}\}$ , $\triangleright$ Sensor info in the i-th frame
6:$v=(id,x,y,z,h,w,l,\theta)$triangleright BBox parameters of the j-th vehicle
7:$s=(id,x,y,z,\theta)$triangleright Sensor parameters of the j-th vehicle
8:Output: Generated dataset $\mathcal{D}$
9:Initialize: $\mathcal{G}=\emptyset$, $\mathcal{D}=\emptyset$, $\delta t=\Delta t/10$
10:
11:for $i\leftarrow 2$ to $N_{f}$ do $\triangleright$ Loop over all frames
12:     for $k\leftarrow 1$ to $10$ do $\triangleright$ Interpolate adjacent two frames into 10 subframes
13:         $\mathbf{v}=\mathbf{s}=\emptyset$
14:         for $j\leftarrow 1$ to $N_{v}$ do $\triangleright$ Interpolate over the parameters of each vehicle
15:              if $v_{i-1,j}$ exists then
16:                  $\mathbf{v}\leftarrow\mathbf{v}+\text{Interpolate}(v_{i-1,j},v_{ij},n$)
17:              end if
18:         end for
19:         for $j\leftarrow 1$ to $N_{s}$ do $\triangleright$ Interpolate over the parameters of each vehicle
20:              if $s_{i-1,j}$ exists then
21:                  $\mathbf{s}\leftarrow\mathbf{s}+\text{Interpolate}(s_{i-1,j},s_{ij},n$)
22:              end if
23:         end for
24:         $t=t_{i}+k\cdot\delta t$triangleright New timestamp for current subframe
25:         $\mathcal{G}\leftarrow\mathcal{G}+(\mathbf{v},\mathbf{s},t)$triangleright Save data for current subframe
26:     end for
27:end for
28:$\mathcal{P}=\emptyset$
29:for $(\mathbf{v},\mathbf{s},t)$ in $\mathcal{G}$ do $\triangleright$ Simulation replay over all new subframes
30:     $\mathbf{p}\leftarrow$ CarlaPlay$(\mathbf{v},\mathbf{s},t)$triangleright Generate $1/10$ lidar scans for each vehicle in $g$
31:     $\mathcal{P}\leftarrow\mathcal{P}+(\mathbf{p},t)$triangleright Save scanned point clouds
32:end for
33:$\mathcal{P}\leftarrow\text{Sort}_{t}(\mathcal{P})$triangleright Sort data in ascending order of $t$
34:$t_{\text{max}}=\max({t|(\mathbf{p},t)\in\mathcal{P}})$
35:for $j\leftarrow 1$ to $N_{s}$ do $\triangleright$ Compose full scans with random sensor ticking times
36:     $\tau=\text{RandomInteger}(1,5)\cdot\delta t$triangleright Generate random sensor ticking time
37:     while $\tau<t_{\text{max}}$ do
38:         $P\leftarrow\{p_{j}|(p_{j},s_{j})\in\mathbf{p},(\mathbf{p},t)\in\mathcal{P},\tau\leq t<\tau+\Delta t\}$triangleright Full scan
39:         $s=s_{j}$ where $(p_{j},s_{j})\in\mathbf{p},(\mathbf{p},t)\in\mathcal{P},t=\tau$triangleright Sensor pose
40:         $G\leftarrow\mathbf{v}\text{ where }(\mathbf{v},\mathbf{s},\tau+\Delta t)\in\mathcal{G}$triangleright Ground-truth BBoxes
41:         $\mathcal{D}\leftarrow\mathcal{D}+(P,G,s,\tau)$
42:         $\tau\leftarrow\tau+\Delta t$
43:     end while
44:end for
Appendix B: DairV2Xt Generation
The overall process of generating DairV2Xt dataset is described in Algorithm 2. Note that this algorithm is applied to each sequence containing $N_{f}$ frames. The functions used in Algorithm 2 are defined as follows:
Register$(O,R)$: Given the reference point cloud $O$ and the source point cloud $R$, this function registers $R$ to $O$. The Iterative Closest Point (ICP) algorithm is used for the registration with a threshold distance of $0.2$m. Note that during registration, dynamic points are moved via the BBox annotations. The registration result is a transformation matrix $M$, which is then used to align the point cloud $R$ to $O$. The transformed $R$ is merged into $O$, and the merged point cloud is downsampled with a voxel size of $0.1$m to speed up the registration process. The Register function returns the updated reference point cloud $O$ and the sensor parameters $s$ obtained from $M$.
Transform$(\mathbf{v},s)$: This function constructs a transformation matrix from $s$ and then transforms the parameters of each $v\in\mathbf{v}$ using the matrix. It returns the transformed bounding box (BBox) parameters.
RetriveTimestamp(v, R): Given the BBox parameters $v$ and the point cloud $R$, this function finds the points $P=\{p\mid p\in R\}$ that lie within the BBox $v$ and returns the mean timestamp of the points in $P$.
Track$(\mathcal{T},\mathbf{u})$: The input $\mathcal{T}$ contains $N$ trajectories, where each trajectory $T_{i}$ is a list of BBoxes $[v_{1},v_{2},\dots,v_{j},\dots,v_{n}]$ with $v_{j}=((x,y,z,h,w,l,\theta),t)$. Each input BBox $u_{k}\in\mathbf{u}$ has the same parameterization as $v_{j}$. For each $u_{k}$, the Euclidean distance $d_{i}$ is calculated between $u_{k}$ and each $v_{n}\in\mathcal{T}$. If $d_{i}<3$m, $u_{k}$ is appended to $T_{i}$; otherwise, a new trajectory $T_{N+1}=[u_{k}]$ is initialized and added to $\mathcal{T}$. The updated tracklets $\mathcal{T}$ are then returned.
Interpolate$(\mathcal{T},t)$: For each trajectory $T_{i}$ in $\mathcal{T}$, the function finds the two BBoxes $b1=(v1,t1)$ and $b2=(v2,t2)$ that are closest in time and fulfill the condition $t1<t\leq t2$. If such BBoxes exist, interpolation is performed using
$u=u1+(u2-u1)\cdot\frac{t-t1}{t2-t1}$ (7)
for the location $x,y,z$ and the orientation $\theta$ of the BBox. The mean dimensions $h,w,l$ of the two BBoxes $v1$ and $v2$ are used for the new interpolated BBox.
Algorithm 2 DairV2Xt Generation
Input: DairV2X dataset $\mathcal{F}$
2:$\Delta t$triangleright Time interval between two adjacent frames
=$\mathcal{F}=\{(P_{i},Q_{i},\mathbf{v}^{P}_{i},\mathbf{v}^{Q}_{i},s^{P}_{i},s^{Q}_{i},t_{i})|i\in\{1,2,\dots,N_{f}\}\}$, $\triangleright$ Frame data at all times $t_{i}$
4:$P=\{x_{i},y_{i},z_{i},t_{i}|i\in\{1,2,\dots,N_{P}\}\}$, $\triangleright$ CAV Point cloud in each frame
$Q=\{x_{i},y_{i},z_{i},t_{i}|i\in\{1,2,\dots,N_{Q}\}\}$, $\triangleright$ Infrastructure Point cloud in each frame
6:$\mathbf{v}^{*}=\{v_{i}|i=\{1,2,\dots,N_{v}\}\}$ , $\triangleright$ BBoxes in point cloud *
$s^{*}=(x,y,z,\theta)$triangleright Sensor parameters for point cloud *
8:$v=(x,y,z,h,w,l,\theta)$triangleright BBox parameters
Output: Generated dataset $\mathcal{D}$
10:Initialize: $\mathcal{D}=\emptyset$
12:$idx\leftarrow\operatorname*{arg\,min}_{i}dist(s^{P}_{i},s^{Q}_{i})$triangleright Index of min. norm-2 distance of sensor positions
$O_{\text{ref}}\leftarrow P_{idx}$triangleright Global reference frame
14:for $i\leftarrow idx$ to $N_{f}$ do $\triangleright$ Register $P$ for $i>idx$
     $(O_{\text{ref}},s^{P}_{i})\leftarrow\text{Register}(O_{\text{ref}},P_{i})$
16:end for
for $i\leftarrow idx$ to $1$ do $\triangleright$ Register $P$ for $i<idx$
18:     $(O_{\text{ref}},s^{P}_{i})\leftarrow\text{Register}(O_{\text{ref}},P_{i})$
end for
20:for $i\leftarrow 1$ to $N_{f}$ do
     $(\cdot,s^{Q}_{i})\leftarrow\text{Register}(O_{\text{ref}},Q_{i})$triangleright Register $Q$
22:     $\mathbf{v}^{P}_{i}\leftarrow\text{Transform}(\mathbf{v}^{P}_{i},s^{P}_{i})$triangleright Transform BBoxes of $P$ to aligned frame $O_{\text{ref}}$
     $\mathbf{v}^{Q}_{i}\leftarrow\text{Transform}(\mathbf{v}^{Q}_{i},s^{Q}_{i})$triangleright Transform BBoxes of $Q$ to aligned frame $O_{\text{ref}}$
24:     for $(\mathbf{v},R_{i})\text{ in }\{(\mathbf{v}^{P}_{i},P_{i}),(\mathbf{v}^{Q}_{i},Q_{i})\}$  do
         $\mathbf{u}\leftarrow\emptyset$
26:         for $v$ in $\mathbf{v}$ do $\triangleright$ Timestamping each BBox
              $t\leftarrow\text{RetriveTimestamp}(v,R_{i})$
28:              $\mathbf{u}\leftarrow\mathbf{u}+(t,v)$
         end for
30:         $\mathcal{T}\leftarrow\text{Track}(\mathcal{T},\mathbf{u})$triangleright Track BBoxes
     end for
32:end for
for $i\leftarrow 1$ to $N_{f}$ do
34:     $t\leftarrow\max_{t}(\{t|(x,y,z,t)\in P_{i}\})$triangleright Align global timestamp to maximum $t$ in $P$
     $\mathbf{v}\leftarrow\text{Interpolate}(\mathcal{T},t)$triangleright Get BBoxes at time $t$
36:     $\mathcal{D}\leftarrow\mathcal{D}+(P_{i},Q_{i},s^{P}_{i},s^{Q}_{i},\mathbf{v})$triangleright Save data for frame $i$
end for